# Multi-Label Bayesian Active Learning with Inter-Label Relationships

### v1

@@ 5 Conclusion
In this paper, we proposed an innovative MLAL query strategy, CRAB, which takes into account inherent label relationships within a Bayesian framework. By updating the correlation matrices with the annotated data, our model is competent to query more representative samples in the initial stage and achieves a more accurate score for evaluating the informativeness of instances. Additionally, with the utilization of beta scoring rules, our model maintains consistently robust performance on imbalanced datasets. Leveraging pseudo labels and correlation-aware sampling, our strategy eliminates the need for additional training modules, and our model demonstrates significant performance improvements in MLAL on four benchmark datasets.
6 Limitation
Our work has two main limitations. First, although our model utilizes subsets of the unlabeled pool to select samples based on correlation properties, the framework could be improved by dynamically and proportionally adjusting the subset sample size for the refined unlabeled pool. Since learning performance varies across datasets, dynamically adapting the sample selection process in response to the model’s evolving learning capability and the inter-relationships within each dataset could further enhance performance, allowing the model to balance diverse and informative sample selection more effectively.
Second, our current approach focuses primarily on label-wise correlations, considering both the co-occurrence and non-co-occurrence relationships between labels. Future work will extend this to explore correlations in both instance and label spaces, examining whether the alignment between instances and labels can further improve performance. Additionally, we plan to investigate more complex relationships, such as spurious correlations, that may exist between data features and label distributions. This includes studying the impact of invalid or noisy information from instances on model performance and identifying methods to mitigate such effects.

### v2

@@ 5 Conclusion
In this paper, we proposed an innovative MLAL query strategy, CRAB, which takes into account inherent label relationships within a Bayesian framework. By updating the correlation matrices with the annotated data, our model is competent to query more representative samples in the initial stage and achieves a more accurate score for evaluating the informativeness of instances. Additionally, with the utilization of beta scoring rules, our model maintains consistently robust performance on imbalanced datasets. Leveraging pseudo labels and correlation-aware sampling, our strategy eliminates the need for additional training modules, and our model demonstrates significant performance improvements in MLAL on four benchmark datasets. Future research could explore the correlation at the instance space and investigate additional relationships between the data features and label distributions.
Acknowledgements.
We would like to thank the anonymous reviewers for their valuable and helpful comments. This research was supported by an Australian Government Research Training Program (RTP) Scholarship.